OpenStudy (anonymous):

A Chemist has a 20-mg sample of polonium-218. He needed approximately 81.5% of it for an experiment. given that the half-life of polonium-218 is approximately 3.1 min, how many seconds will it take for the sample to decay to the desired mass

OpenStudy (anonymous):

52 seconds

OpenStudy (aaronq):

since it's a first-order process, you can use the integrated rate law equation: \(ln[A]_t=-kt+ln[A]_o\) where \([A]_t\) is the amount of substance after t time has elapsed. \([A]_o\) is initial amount k is the decay constant t is time you can sub in: \(t_{1/2}=\dfrac{ln(2)}{k}\); where \(t_{1/2}\) is the half-life you end up with: \(ln[A]_t=-\dfrac{ln(2)*t}{t_{1/2}}+ln[A]_o\) plug and chug. The answer is 49.64 seconds

OpenStudy (anonymous):

i knew that